If P(A) = 0.80, P(B) = 0.70 AND P(A or B) = 0.90, then P(A and B) is what?
Answer by robertb(5830)

P(A and B) = P(A) + P(B) - P(A or B) = 0.80 + 0.70 -0.90 =0.6